OpenStudy (anonymous):

For the graph shown, select the statement that best represents the given system of equations. algebra unit 8 lesson 1 assessment question 2 2y – x = 2 x – 2y = 8 A. consistent and independent B. not enough information C. inconsistent D. coincident

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

plz help

OpenStudy (alexandervonhumboldt2):

one second

OpenStudy (anonymous):

ok

OpenStudy (alexandervonhumboldt2):

If the two lines are parallel, the system is inconsistent, with no solution.

OpenStudy (alexandervonhumboldt2):

C

OpenStudy (anonymous):

thanks can you help with more plz?

OpenStudy (alexandervonhumboldt2):

yes

OpenStudy (anonymous):

For the graph shown, select the statement that best represents the given system of equations. 3y – x = 12 x – 3y = 6 A. coincident B. consistent and independent C. inconsistent D. not enough information

OpenStudy (anonymous):

C?

OpenStudy (anonymous):

ok

OpenStudy (alexandervonhumboldt2):

they are parallel so C

OpenStudy (anonymous):

How many solutions does this system of equations have? A. none B. exactly one C. exactly two D. infinitely many

OpenStudy (anonymous):

confused on this one

OpenStudy (anonymous):

A right?

OpenStudy (alexandervonhumboldt2):

A

OpenStudy (alexandervonhumboldt2):

they dont intersect so none

OpenStudy (anonymous):

The system of equations is inconsistent. What are the missing values?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

give me a sec i have to writw ir out

OpenStudy (alexandervonhumboldt2):

ok

OpenStudy (anonymous):

The system of equations is inconsistent. What are the missing values? y=2/3x+3 y=____/____x+5

OpenStudy (alexandervonhumboldt2):

2/3

OpenStudy (alexandervonhumboldt2):

other questions plase post in separate post ans tag me

OpenStudy (anonymous):

ok so the answer is 2/3 for the blank?

OpenStudy (alexandervonhumboldt2):

yes

OpenStudy (anonymous):

ok give me a sec

OpenStudy (alexandervonhumboldt2):

pralle lines have the same slope as incostinent have are parallel then 2/3

OpenStudy (anonymous):

ok thx